OpenStudy (ryuga):

Which expression is equivalent to this one? u • 5 + u • v A. u(5 + v) B. 5(u + v) C. v(5 + u) D. u · (5 + u) · v

OpenStudy (tkhunny):

Let's try the Distributive Property of Multiplication over Addition, shall we?

OpenStudy (ryuga):

ok?

OpenStudy (ryuga):

@skullpatrol

OpenStudy (ryuga):

@tkhunny

OpenStudy (ryuga):

@SamsungFanBoy

OpenStudy (ryuga):

@dakid88

OpenStudy (ryuga):

@mathstudent55

OpenStudy (dakid88):

hold on one sec

OpenStudy (tkhunny):

So, your not going to try it?

OpenStudy (ryuga):

well what am i supposed to do?

OpenStudy (dakid88):

C.

OpenStudy (tkhunny):

Are you saying you've never heard of the distributive property? This one: a(b+c) = ab + ac

OpenStudy (ryuga):

no

OpenStudy (ryuga):

i know a+b=b+a

OpenStudy (tkhunny):

That's the commutative property. Good. a(bc) = (ab)c -- is associative. That also works for addition.

OpenStudy (tkhunny):

BTW - C is not correct. Not sure what dakid88 was doing. I got one gloriously wrong, yesterday. It happens.

OpenStudy (tkhunny):

So, apply this principle a(b+c) = ab + ac to your problem u • 5 + u • v What do you get?